Please give me the answer for


Insert a A.M. betweeb 5 and 20.


 


Anitha.


 

Grade:Upto college level

2 Answers

vikas askiitian expert
13 years ago

AM(arithemetic mean) = a+b/2

here in this question a = 5 & b = 20 so

AM = 20+5/2 = 25/2

now series is 5 , 25/2 , 20

common difference = 15/2 & first term is 5

 

approve if u like my ans

Sudheesh Singanamalla
13 years ago

assume the arithmetic mean to be A

so the sequence now is

5 , A , 20

A = (a+b)/2 where a = 5 , b = 20

so A = ( 20+5 )/2

A = 25/2

 

Please approve